Erie Company manufactures a mobile fitness device called the Jogging Mate. The company uses standards to control its costs. The labor standards that have been set for one Jogging Mate are as follows:

Standard
Hours
Standard Rate
per Hour
Standard
Cost
27 minutes $6.40 $2.88

During August, 9,470 hours of direct labor time were needed to make 19,800 units of the Jogging Mate. The direct labor cost totaled $58,714 for the month.

Required:

1. What is the standard labor-hours allowed (SH) to makes 19,800 Jogging Mates?

2. What is the standard labor cost allowed (SH × SR) to make 19,800 Jogging Mates?

3. What is the labor spending variance?

4. What is the labor rate variance and the labor efficiency variance?

5. The budgeted variable manufacturing overhead rate is $4.50 per direct labor-hour. During August, the company incurred $47,350 in variable manufacturing overhead cost. Compute the variable overhead rate and efficiency variances for the month.

(For requirements 3 through 5, indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance). Input all amounts as positive values. Do not round intermediate calculations.)
Answer #1

1. Standard labor hours allowed to make 19,800 Jogging Mates = 27 / 60 x 19,800 = 8,910 hours.

2. Standard labor cost allowed to make 19,800 Jogging Mates = 19,800 x $ 2.88 = $ 57,024

3. Labor spending variance = $ 58,714 - $ 57,024 = $ 1,690 U

4. Labor rate variance = $ ( 6.20 - 6.40 ) x 9,470 hours = $ 1,894 F

Labor efficiency variance = $ ( 9,470 - 8,910 ) x $ 6.40 = $ 3,584 U

5. Variable overhead rate variance = $ ( 5.00 - 4.50 ) x 9,470 hours = $ 4,735 U

Variable overhead efficiency variance = ( 9,470 - 8,910 ) hours x $ 4.50 = $ 2,520 U